Help with managerial accounting??

Karney Company has revenues of $500,000, variable costs of $350,000, and fixed costs of $135,000. The sales price per unit is $100 and the Variable Cost per unit is $70, at this level of sales the Fixed Costs per unit is $27.

a. compute the contribution margin per unit and the contribution margin ratio

b. compute total unfits and sales dollars needed to break even.

c. compute total sales dollars needed to achieve a target operating income of $45,000
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

H А в DE Answer Part 1) Contribution Margin Per Unit = sale price - variable cost = 100-70 = $ 30 Contribution Margin ratio =

7 Part 2) B Break even units = Fixed Cost/contribution Margin per unit 9 = 135000/30 0 = 4500 Units 2 Break even sales dollar